What is the y=mx+b form of "x-4y=12"
Found 2 solutions by nerdybill, rfer:
Answer by nerdybill(7384)

What is the y=mx+b form of "x-4y=12"
.
starting with:
x-4y=12
-4y=-x+12
4y=x-12
y=(1/4)x-3

Answer by rfer(16322)

-4y=-x+12
y=1/4x-3